# Meta-Task Prompting Elicits Embedding from Large Language Models

### v1

@@ Abstract
In this work, we introduce a new unsupervised embedding method, Meta-Task Prompting with Explicit One-Word Limitation (MetaEOL), for generating high-quality sentence embeddings from Large Language Models (LLMs) without the need for model fine-tuning or task-specific engineering. Leveraging meta-task prompting, MetaEOL guides LLMs to produce embeddings through a series of carefully designed prompts that address multiple representational aspects. Our comprehensive experiments demonstrate that embeddings averaged from various meta-tasks yield competitive performance on Semantic Textual Similarity (STS) benchmarks and excel in downstream tasks, surpassing contrastive-trained models. Our findings suggest a new scaling law for embedding generation, offering a versatile, resource-efficient approach for embedding extraction across diverse sentence-centric scenarios.
Meta-Task Prompting Elicits Embedding from Large Language Models
Yibin Lei1, Di Wu1, Tianyi Zhou2, Tao Shen3, Yu Cao4, Chongyang Tao5, Andrew Yates1 1University of Amsterdam  2University of Maryland 3University of Technology Sydney  4Tecent IEG  5Microsoft Corporation {y.lei, d.wu, a.c.yates}@uva.nl, tianyi@umd.edu tao.shen@uts.edu.au, rainyucao@tencent.com, chotao@microsoft.com
@@ 1 Introduction
The advent of Large Language Models (LLMs) such as GPT-3 (Brown et al., 2020) and LLaMA (Touvron et al., 2023a) has marked a significant milestone in the field of natural language processing (NLP), introducing promising unsupervised methods for various NLP tasks by leveraging task-related instructions or prompts. These tasks also include the generation of sentence embeddings, which aims to produce sentence representations that can be applied across a spectrum of scenarios, ranging from intrinsic tasks like Semantic Textual Similarity (STS) (Agirre et al., 2012a; Cer et al., 2017b) to downstream tasks including information retrieval (Mitra et al., 2017; Izacard et al., 2021), sentiment categorization (Ke et al., 2020), and beyond. By employing specific prompts (Jiang et al., 2023b, 2022a), researchers have begun to explore the potential of extracting meaningful sentence embeddings directly from the hidden states of LLMs without the need for explicit training on embedding-specific tasks. The significance of zero-resource settings – where embeddings are generated without any fine-tuning or in-context learning – cannot be overstated.
@@
Figure 1: The highest decoding probabilities are largely allocated to stop words that carry little useful information when conducting a meaning compression prompting, even if employing a constraint of "in one word" following  (Jiang et al., 2023b). Although the general semantic, movie, is contained, other aspects of this sentence are missing, like sentiments.
@@
Initial efforts in this domain, as highlighted by works like (Jiang et al., 2023b, 2022a; Liu et al., 2023a), have focused on unsupervised techniques that extract sentence representations directly from LLMs. These methods typically involve using fill-in-the-blanks prompts, such as This sentence: "[TEXT]" means in one word:" (Jiang et al., 2023b), to embed a sentence into a single token representation, achieved by extracting output hidden states of the last token as the sentence’s embedding. Despite promising, they also reveal the inherent challenges where embeddings may be overly simplistic or misaligned with the intended semantic nuances of the sentences.
@@
In our pilot experiment, we show that the previous prompt-based method (Jiang et al., 2023b) struggles to capture a comprehensive meaning for a sentence, especially when the usage of the sentence is associated with multiple aspects. As illustrated in Figure 1, when probing probability distribution for the next token during decoding, which reflects the embedding quality of the last token, the highest probabilities are mostly distributed to frequent stop words. Although the general semantic — movie — is contained, other meaningful aspects like sentiments are missing.
@@
A straightforward solution to mitigate this issue is to demonstrate LLMs with task-specific instructions. This approach involves instructing the model with prompts explicitly designed for a particular task, thereby tailoring the embeddings to better suit the specific requirements of that task. However, considering the realm of NLP encompasses thousands of distinct tasks (Mishra et al., 2022; Wang et al., 2022; Chung et al., 2022), this would be impractical due to computational and storage challenges. Furthermore, task-specific embeddings, while effective for their designated tasks, often fail to generalize well across different tasks.
@@
Drawing on the principles of the usage-based theory of language acquisition (Tomasello, 2009), which asserts that the essence of meaning is rooted in the practical utilization of language, we aim to introduce a novel, unsupervised approach for the extraction of high-quality sentence embeddings directly from LLMs. Our method does not confine itself to deriving embeddings from a singular, overarching meaning. Instead, it embraces the concept of meta-task prompting, inspired by the burgeoning fields of meta-task prompted training (Sanh et al., 2022) and hyper-prompt (He et al., 2022) techniques. By defining a suite of meta-tasks, each tailored to a distinct application context, MetaEOL prompts LLMs to consider multiple representational tokens from a variety of perspectives, as shown in Figure 2. This multifaceted approach enables the extraction of more diverse and nuanced contextualized embeddings that collectively form a comprehensive sentence embedding.
@@
Extensive experiments empirically show that: (I) Simply averaging embeddings from different meta-tasks without any training leads to general embeddings that are competitive to contrastive-trained models on STS tasks and can achieve the best average result in downstream tasks. (II) Incrementally integrating more meta-tasks (ranging from one to four) yields consistent improvements across STS tasks, showcasing high generalities, and highlighting the significant impact of meta-task integration on overall performance. (III) The final layer is not always the most effective for STS tasks and with a simple proportional layer selection strategy, we achieve the best results with a 70B model, which points to a potential scaling law.
@@
Figure 2: The workflow of our method (MetaEOL). We use the prompt in Appendix A.1 to prompt ChatGPT-4 to generate templates. Each input sentence will be decorated with multiple task-specific templates, indicating its various intended usage scenarios. The resulting multiple prompts will be fed to LLMs. Then, multiple task-specific embeddings will be extracted.
@@ 2 Related Work
Sentence embeddings aim to encapsulate the semantic content of sentences into fixed-sized vector representations. Recent developments in contrastive learning have proven to be highly effective for generating sentence embeddings, under both unsupervised and supervised settings (Gao et al., 2021; Jiang et al., 2022a; Chuang et al., 2022; Wu et al., 2022). For instance, SimCSE (Gao et al., 2021) utilizes different dropout masks as a form of noise to create positive pairs in an unsupervised fashion, while in a supervised setting, models like Sentence-BERT Reimers and Gurevych (2019) leverage natural language inference (NLI) datasets to construct positive and negative pairs. Additionally, Su et al. (2023) and Asai et al. (2023) show that training with a large amount of tasks with annotated instructions can enable the model to generate embeddings tailored to different downstream tasks. In contrast, our approach MetaEOL demonstrates the potential of utilizing LLMs directly to generate instruction-followed embeddings without necessitating any training.
Large Language Models for Sentence Representation.
Recent studies have explored the application of LLMs for enhancing sentence embeddings through data augmentation techniques (Cheng et al., 2023; Zhang et al., 2023; Xu et al., 2024). Notably, Sentence-T5 Ni et al. (2022) employs contrastive learning on models with billions of parameters, demonstrating consistent performance improvements correlating with increases in model size. More recently, Liu et al. (2023b) represents sentences through the distribution of possible text continuations, comparing the distributional similarity between sentences. This method, although effective, necessitates the generation of 20 trajectories, each up to 20 tokens in length, making it computationally intensive. Moreover, this type of representation can not directly be used for downstream tasks. Jiang et al. (2022a) incorporates in-context learning (Dong et al., 2023) to enhance sentence embeddings. While proven effective, it also reveals that the produced embeddings are task-specific and struggle with generalization across various downstream tasks, in addition to being highly sensitive to the choice of demonstrations.
@@ 3 Method
In this section, we begin by reviewing two kinds of previous prompting methods for deriving sentence representation from masked and causal language models, respectively (see Section 3.1). Subsequently, we describe our proposed method, i.e., Meta-Task Prompting with Explicit One-Word Limitation (MetaEOL) (see Section 3.2) in detail. Lastly, we describe meta-tasks involved in this paper (see Section 3.3).
@@ 3.1.2 Causal Language Model
There are trials to directly extract sentence representation from large Causal Language Models (CLMs), e.g., OPT Zhang et al. (2022) or LLAMA Touvron et al. (2023a), without additional training. Inspired by  (Jiang et al., 2022a), PromptEOL (Jiang et al., 2023b) employs a similar template as follows,
@@
for prompting, where the last layer’s hidden vector for the last token “” is extracted as the sentence representation. A constraint of “in one word” is applied to avoid the model’s tendency to generate long sentences such that the last token fails to capture the overall information.
Nevertheless, the obtained embedding highly relies on the single prompt. So the inference process is confined and results in incomprehensive features, making it not applicable to extensive scenarios. As shown in Figure 1, for a negative review of a movie, the resulting embedding only contains the general semantics of movies but falls short of capturing other critical aspects such as sentiment.
@@ 3.2 Meta-Task Prompting
To overcome the issues raised above, we propose Meta-Task Prompting with Explicit One-Word Limitation (MetaEOL). A meta-task is associated with a potential broad usage scenario for the corresponding sentence representation. As shown in Figure 2, we directly prompt causal LLMs with the goals of multiple meta tasks, aiming to obtain the representations under various broad intents.
Specifically, we produce task-oriented prompts by decorating the original prompting template used for causal LLMs (see Section 3.1.2) with the corresponding task description. For example, given a meta-task where representations are extracted for Text Classification (TC), we extend the template with task-oriented context to define the behavior during inference. As the template of Meta Task-1 shown in Figure 2, a detailed task description text, telling the LLM that it should categorize the excerpt into a broad category, is placed at the beginning of the prompt. Then, an instruction with a constraint of "in one word" is followed to ensure models aggregate the information of the whole sentence into the embedding of the last token. The placeholder [TEXT] will be substituted with the original sentence to produce the final task-oriented prompt. The resulting task-specific prompt will serve as input to LLMs. Subsequently, we extract the hidden vector of the last token “” as the sentence representation, following the pattern outlined in Section 3.1.
It is worth noting that given various meta-tasks, distinct templates will be employed, leading to multiple different sentence embeddings. Our hypothesis posits that each embedding captures a distinct representation customized for a specific feature viewpoint (meta-task). In this paper, we empirically show that simply averaging different embedding derived from multiple meta-tasks can achieve superior performance for both intrinsic and downstream evaluation benchmarks.
@@ 3.3 Types of Meta-Tasks
In this paper, we conduct experiments on the following four distinct meta-tasks, i.e., Text Classification (TC), Sentiment Analysis (SA), Paraphrase Identification (PI), and Information Extraction (IE), aiming to capture information from different angles. E.g., intuitively, the TC task primarily emphasizes topic-level information, whereas the IE task concentrates on surface-level signals.
@@
Note that introducing more meta-tasks is trivial, which only requires adding more task names to the generator. Here, we choose the above four meta-tasks as a testbed to assess the scalability. More specifically, in Section 5.2, we show that incrementally adding more meta-tasks to our workflow results in consistently better performance.
@@ 4 Experiments
PromptEOL-LLAMA2 13B 56.19 76.42 65.42 7273 75.21 67.96 68.23 68.83
@@ 4.1 Settings
Suggested by prior works (Reimers and Gurevych, 2019; Gao et al., 2021; Jiang et al., 2022b) that the primary objective of sentence embeddings is to cluster semantically similar sentences, we evaluate MetaEOL on seven semantic textual similarity (STS) datasets, utilizing the SentEval toolkit (Conneau and Kiela, 2018). The STS datasets consist of STS 2012-2016 (Agirre et al., 2012b, 2013, 2014, 2015, 2016), STS-B (Cer et al., 2017a), and SICK-R Marelli et al. (2014). Each sentence pair in the STS datasets is annotated with a score from 0 to 5 indicating the pairwise semantic similarity. The Spearman correlation (scaled by 100x) between the model-predicted similarity scores and human-annotated similarity scores is used as the metric. We employ cosine similarity to measure the similarity between sentence embeddings. The Spearman correlation is computed under the “all” setting.
@@
The baselines we consider can be mainly categorized into two types – models with contrastive training and without contrastive training: (I) Models with Contrastive Training: We compare MetaEOL with SOTA unsupervised contrastive-trained models, namely SimCSE (Gao et al., 2021) and PromptBERT Jiang et al. (2022a). The models are trained on $10^{6}$ sentences randomly sampled from Wikipedia. Results based on BERT (Devlin et al., 2019) and RoBERTa Liu et al. (2019) models are reported. And, (II) Models without Contrastive Training: We compare MetaEOL with (1) Average pooling methods, where average pooling is applied to the output hidden states of all tokens in a sentence to obtain the sentence embedding. We report results with BERT, the encoder of ST5 (Ni et al., 2022), LLAMA2 (Touvron et al., 2023b) and Mistral (Jiang et al., 2023a) models; and (2) Prompt-based methods, which include BERT Prompt that employs the same prompt strategy as PromptBERT but does not incorporate contrastive training, and also PromptEOL. All methods mentioned above rely on the output from the final layer to obtain the sentence embedding.
Implementation Details.
We apply MetaEOL to LLaMA2-7B, LLAMA2-13B, and Mistral-7B models, using meta-tasks consisting of Text Classification (TC), Sentiment Analysis (SA), Paraphrase Identification (PI), and Information Extraction (IE). These tasks are distinct and collectively consider diverse aspects of a sentence. For each of these meta-tasks, we utilize GPT-4 to create two unique task prompts, resulting in a total of eight task prompts. [^1] The details of these eight task prompts are presented in Appendix A.3. MetaEOL rely on the output from the final layer to obtain the sentence embedding. We simply average the resulting embeddings of task prompts from different meta-tasks to obtain the final embedding.

### v2

# Meta-Task Prompting Elicits Embeddings from Large Language Models

@@ Abstract
We introduce a new unsupervised text embedding method, Meta-Task Prompting with Explicit One-Word Limitation (MetaEOL), for generating high-quality sentence embeddings from Large Language Models (LLMs) without the need for model fine-tuning. Leveraging meta-task prompting, MetaEOL guides LLMs to produce embeddings through a series of carefully designed prompts that address multiple representational aspects. Our comprehensive experiments demonstrate that embeddings averaged from various meta-tasks are versatile embeddings that yield competitive performance on Semantic Textual Similarity (STS) benchmarks and excel in downstream tasks, surpassing contrastive-trained models. Our findings suggest a new scaling law, offering a versatile and resource-efficient approach for embedding generation across diverse scenarios. [^1] Our code is publicly available at https://github.com/Yibin-Lei/MetaEOL.
Meta-Task Prompting Elicits Embeddings from Large Language Models
Yibin Lei1thanks: Corresponding to: Yibin Lei (e-mail: y.lei@uva.nl) and Chongyang Tao (e-mail: chotao@microsoft.com)., Di Wu1, Tianyi Zhou2, Tao Shen3, Yu Cao4, Chongyang Tao511footnotemark: 1, Andrew Yates1 1University of Amsterdam  2University of Maryland 3AAII, FEIT, University of Technology Sydney  4Tencent IEG  5Microsoft Corporation {y.lei, d.wu, a.c.yates}@uva.nl, tianyi@umd.edu tao.shen@uts.edu.au, rainyucao@tencent.com, chotao@microsoft.com
@@ 1 Introduction
The advent of Large Language Models (LLMs) such as GPT-3 (Brown et al., 2020) and LLaMA (Touvron et al., 2023a) has marked a significant milestone in the field of natural language processing (NLP), introducing promising unsupervised methods for various NLP tasks by leveraging task-related instructions or prompts (Qin et al., 2023; Zhong et al., 2023; Zhao et al., 2023). These tasks also include the generation of sentence embeddings, which aims to produce sentence representations that can be applied across a wide range of scenarios. They have been applied to intrinsic tasks like Semantic Textual Similarity (STS) (Agirre et al., 2012a; Cer et al., 2017b), to downstream tasks including information retrieval (Mitra et al., 2017; Izacard et al., 2021), and to sentiment classification (Ke et al., 2020) and beyond. By employing specific prompts (Jiang et al., 2023b, 2022a), researchers have begun to explore the potential of extracting meaningful sentence embeddings directly from the hidden states of LLMs without the need for explicit training. These prompt-based approaches generate embeddings without the need for any fine-tuning or in-context learning, which is a substantial improvement over approaches that require extensive fine-tuning to achieve high performance.
@@
Figure 1: The highest decoding probabilities are largely allocated to stop words that carry little useful information when conducting a meaning compression prompting, even if employing a constraint of "in one word" following Jiang et al. (2023b). Although the general semantic, movie, is contained, other aspects of this sentence are missing, like sentiments.
@@
Initial efforts in this domain, as highlighted by works like (Jiang et al., 2023b, 2022a; Liu et al., 2023a), have focused on unsupervised techniques that extract sentence representations directly from LLMs. These methods typically involve using fill-in-the-blanks prompts, such as This sentence: "[TEXT]" means in one word:" (Jiang et al., 2023b), to embed a sentence into a single token representation by using the output hidden state of the last token as the sentence’s embedding. While they perform well, these approaches also reveal the inherent challenges of this task: embeddings may be overly simplistic or misaligned with the intended semantic nuances of the sentences.
@@
In a pilot experiment illustrated in Figure 1, we demonstrate that a previous prompt-based method (Jiang et al., 2023b) can struggle to capture a sentence’s meaning, especially when the usage of the sentence is associated with multiple aspects. When probing the probability distribution for the next token during decoding, which reflects the embedding quality of the last token [^2] The decoding probabilities are derived by comparing the similarity between the output hidden state of the last token and the token embeddings of the whole vocabulary., the highest probabilities are mostly distributed to frequent stop words. Although the general movie topic appears, other meaningful aspects like sentiments are missing.
@@
A straightforward solution to mitigate this issue is to provide LLMs with task-specific instructions. This approach involves instructing the model with prompts explicitly designed for a particular task, thereby tailoring the embeddings to better suit the specific requirements of that task. However, considering the wide range of distinct tasks that an embedding may be used for (Mishra et al., 2022; Wang et al., 2022; Chung et al., 2022), this would be impractical. Furthermore, while task-specific embeddings are effective for their corresponding tasks, they may fail to generalize well across different tasks.
@@
Inspired by the principles of the usage-based theory of language acquisition (Tomasello, 2009), which asserts that the essence of meaning is rooted in the practical utilization of language, our approach aims to generate broad embeddings through the use of meta-task prompting, inspired by meta-task prompted training (Sanh et al., 2022) and hyper-prompt (He et al., 2022) techniques. By defining a suite of meta-tasks, each tailored to a distinct application context, MetaEOL prompts LLMs to consider multiple representational tokens from a variety of perspectives. This multifaceted approach enables the extraction of more diverse and nuanced contextualized token embeddings that collectively form a comprehensive sentence embedding.
@@
Extensive experiments empirically show that: (i) Simply averaging embeddings from different meta-tasks without any training leads to general embeddings that are competitive to contrastive-trained models on STS tasks and can achieve the best average result on several downstream tasks. (ii) Incrementally integrating more meta-tasks (ranging from one to four) yields consistent improvements across STS tasks, showcasing high generalities, and highlighting the significant impact of meta-task integration on overall performance. (iii) The final layer is not always the most effective for STS tasks and with a simple proportional layer selection strategy, we achieve the best results with a 70B model, which points to a potential scaling law.
@@
Figure 2: The workflow of our method (MetaEOL). We use the prompt in Appendix A.1 to prompt ChatGPT-4 to generate templates. Each input sentence will be decorated with multiple task-specific templates, indicating its various intended usage scenarios. The resulting multiple prompts will be fed to LLMs. Then, multiple task-specific embeddings will be extracted. The final sentence embedding is obtained by averaging the task-specific embeddings.
@@ 2 Related Work
Sentence embeddings aim to encode the semantic content of sentences into fixed-sized vector representations. Recent developments in contrastive learning have proven to be highly effective for generating sentence embeddings, under both unsupervised and supervised settings (Gao et al., 2021; Jiang et al., 2022a; Chuang et al., 2022; Wu et al., 2022). For instance, SimCSE (Gao et al., 2021) utilizes different dropout masks as a form of noise to create positive pairs in an unsupervised fashion, while in a supervised setting, models like Sentence-BERT Reimers and Gurevych (2019) leverage natural language inference (NLI) datasets to construct positive and negative pairs. Additionally, Su et al. (2023) and Asai et al. (2023) show that training with a large amount of tasks with annotated instructions can enable the model to generate embeddings tailored to different downstream tasks. In contrast, our approach MetaEOL demonstrates the potential of utilizing LLMs directly to generate instruction-followed embeddings without the need for any additional training.
Large Language Models for Text Representation.
Recent studies have explored the use of LLMs for enhancing text embeddings through data augmentation techniques (Cheng et al., 2023; Zhang et al., 2023). Notably, Sentence-T5 Ni et al. (2022a) and GTR Ni et al. (2022b) employ contrastive learning on models with billions of parameters. More recently, research has focused on converting an LLM directly into a text encoder without any training. Liu et al. (2023b) represents sentences through the distribution of possible text continuations, comparing the distributional similarity between sentences. This method, although effective, necessitates the generation of 20 trajectories, each up to 20 tokens in length, making it computationally intensive. Jiang et al. (2022a) incorporates in-context learning (Dong et al., 2023) to enhance sentence embeddings. While proven effective, it also reveals that the produced embeddings are task-specific and struggle with generalization across various downstream tasks, in addition to being highly sensitive to the choice of demonstrations. Contemporaneous works investigate the potential of LLMs either by repeating the input texts Springer et al. (2024) or enabling bidirectional attention BehnamGhader et al. (2024) to address the lack of backward dependency in LLMs. Additionally, the potential of LLMs under supervised training settings has also been studied in recent works Ma et al. (2023); Li and Li (2024b, a); Wang et al. (2024); Li et al. (2024); Lee et al. (2024).
Multitask Prompts.
Studies have demonstrated that models fine-tuned using multi-task prompts and datasets can serve as general-purpose models with strong capabilities in generalizing to new tasks (Sanh et al., 2022; Wei et al., 2022; Chung et al., 2022; Wang et al., 2022; Mishra et al., 2022). Our approach MetaEOL aligns with this concept, showcasing that multi meta-task prompts can similarly generate general-purpose embeddings, remarkably without necessitating any training.
@@ 3 Method
In this section, we begin by reviewing two kinds of previous prompting methods for deriving sentence representation from masked and causal language models, respectively (Section 3.1). Subsequently, we describe our proposed method, i.e., Meta-Task Prompting with Explicit One-Word Limitation (MetaEOL) in detail (Section 3.2). Lastly, we describe meta-tasks involved in this paper (Section 3.3).
@@ 3.1.2 Causal Language Model
Others have investigated directly extracting sentence representation from large Causal Language Models (CLMs), e.g., OPT Zhang et al. (2022) or LLAMA Touvron et al. (2023a), without additional training. Inspired by  Jiang et al. (2022a), PromptEOL (Jiang et al., 2023b) employs a similar prompting template as follows,
@@
where the last layer’s hidden vector for the last token “” is extracted as the sentence representation. A constraint of “in one word” is applied to avoid the model’s tendency to generate long sentences such that the last token fails to capture the overall information.
However, the obtained embedding highly relies on the single prompt, which confines the inference process and can result in non-comprehensive features. For example, as shown in Figure 1, for a negative review of a movie, the resulting embedding does not capture critical aspects such as sentiment.
@@ 3.2 Meta-Task Prompting
To overcome the issues raised above, we propose Meta-Task Prompting with Explicit One-Word Limitation (MetaEOL). A meta-task is associated with a potential broad usage scenario for the corresponding sentence representation. As shown in Figure 2, we directly prompt casual LLMs with the goals of multiple meta tasks, aiming to obtain the representations under various broad intents.
Specifically, we produce task-oriented prompts by decorating the original prompting template used for causal LLMs (Section 3.1.2) with the corresponding task description. For example, given a meta-task where representations are extracted for Text Classification (TC), we extend the template with task-oriented context to define the behavior during inference. As shown in the template of Meta Task-1 in Figure 2, a detailed task description text is placed at the beginning of the prompt, instructing the LLM to categorize the excerpt into a broad category. Then, an instruction with a constraint of “in one word” is followed to ensure models aggregate the information of the whole sentence into the embedding of the last token. The placeholder [TEXT] will be substituted with the original sentence to produce the final task-oriented prompt. The resulting task-specific prompt will serve as input to LLMs. Subsequently, we extract the hidden vector of the last token “” as the sentence representation, following the pattern outlined in Section 3.1.
It is worth noting that given various meta-tasks, distinct templates will be employed, leading to multiple different sentence embeddings. Our hypothesis is that each embedding captures a distinct representation customized for a specific feature viewpoint (meta-task). In this paper, we empirically show that simply averaging different embedding derived from multiple meta-tasks can achieve superior performance for both intrinsic and downstream evaluation benchmarks.
@@ 3.3 Types of Meta-Tasks
In this paper, we conduct experiments on the following four distinct meta-tasks, i.e., Text Classification (TC), Sentiment Analysis (SA), Paraphrase Identification (PI), and Information Extraction (IE), aiming to capture information from different angles. For example, intuitively, the TC task primarily emphasizes topic-level information, whereas the IE task concentrates on surface-level signals.
@@
Note that introducing more meta-tasks is trivial, because it only requires adding more task names to the generator. Here, we choose the above four meta-tasks as a testbed to assess scalability. More specifically, in Section 5.2, we show that incrementally adding more meta-tasks to our workflow results in consistently better performance.
@@ 4 Experiments
PromptEOL-LLAMA3 8B 60.88 78.57 68.18 76.75 77.16 72.83 68.94 71.90
PromptEOL-LLAMA2 13B 56.19 76.42 65.42 72.73 75.21 67.96 68.23 68.83
@@ 4.1 Settings
Suggested by prior works (Reimers and Gurevych, 2019; Gao et al., 2021; Jiang et al., 2022b) that an important objective of sentence embeddings is to cluster semantically similar sentences, we evaluate MetaEOL on seven semantic textual similarity (STS) datasets, utilizing the SentEval toolkit (Conneau and Kiela, 2018). The STS datasets consist of STS 2012-2016 (Agirre et al., 2012b, 2013, 2014, 2015, 2016), STS-B (Cer et al., 2017a), and SICK-R Marelli et al. (2014). Each sentence pair in the STS datasets is annotated with a score from 0 to 5 indicating the pairwise semantic similarity. The Spearman correlation (scaled by 100x) between the model-predicted similarity scores and human-annotated similarity scores is used as the metric. We employ cosine similarity to measure the similarity between sentence embeddings. The Spearman correlation is computed under the “all” setting.
@@
The baselines we consider can be categorized into two types – models with contrastive training and without contrastive training: (i) Models with Contrastive Training: We compare MetaEOL with SOTA unsupervised contrastive-trained models, namely SimCSE (Gao et al., 2021) and PromptBERT Jiang et al. (2022a). The models are trained on $10^{6}$ sentences randomly sampled from Wikipedia. Results based on BERT (Devlin et al., 2019) and RoBERTa Liu et al. (2019) models are reported. Contemporaneous LLM-based approach LLM2Vec BehnamGhader et al. (2024) is also included for comparison. LLM2Vec comprises three stages: bidirectional attention enabling, masked next token prediction training, and unsupervised contrastive training (similar to SimCSE) to transform an LLM into a text encoder. Considering (ii) Models without Contrastive Training: We compare MetaEOL with (1) average pooling methods, where average pooling is applied to the output hidden states of all tokens in a sentence to obtain the sentence embedding. We report results with BERT, the encoder of ST5 (Ni et al., 2022a), LLAMA2 (Touvron et al., 2023b) and Mistral (Jiang et al., 2023a) models; and (2) Prompt-based methods, which include BERT Prompt that employs the same prompt strategy as PromptBERT but does not incorporate contrastive training, PromptEOL and the contemporaneous Echo embeddings Springer et al. (2024). Echo embeddings repeat the input once and extract embeddings from the second occurrence. All methods mentioned above rely on the output from the final layer to obtain the sentence embedding.
@@
Implementation Details.
We apply MetaEOL to LLAMA2-7B, LLAMA3-8B, LLAMA2-13B, and Mistral-7B models, using meta-tasks consisting of Text Classification (TC), Sentiment Analysis (SA), Paraphrase Identification (PI), and Information Extraction (IE). These tasks are distinct and collectively consider diverse aspects of a sentence. For each of these meta-tasks, we utilize GPT-4 to create two unique task prompts, resulting in a total of eight task prompts. [^3] The details of these eight task prompts are presented in Appendix A.3. MetaEOL rely on the output from the final layer to obtain the sentence embedding. We simply average the resulting embeddings of task prompts from different meta-tasks to obtain the final embedding.